# RadIR: A Scalable Framework for Multi-Grained Medical Image Retrieval via Radiology Report Mining

### v1

@@ 1 Introduction
The objective of this paper is to develop an image retrieval system for medical applications that ranks instances in a retrieval set based on their relevance to a query, which includes a radiology image and an optional text condition indicating the region to focus on, i.e., the name of anatomy. Such a system has broad implications in enhancing clinicians’ ability to identify similar cases, supporting diagnosis and treatment planning, and facilitating medical education and research [2, 15, 14, 5]. Furthermore, in building generalist models, retrieval-augmented generation (RAG) plays a crucial role in reducing hallucinations and supporting case-based reasoning by grounding outputs in retrieved evidence [19].
@@
To address the challenges we propose a novel medical image ranking pipeline by mining the multi-grained annotations from corresponding radiology reports. Specifically, given a certain anatomy structure, we first standardize the paired reports and extract the relative findings. Then, we adopt the text-level similarity ranking of the findings based on well-designed language-wise metrics [21], to, in turn, represent the image similarity ranks regarding this anatomy structure. This pipeline enables the construction of multi-granular similarity ranking training data in a scalable and automated manner, for both global image matching and fine-grained retrieval conditioned on anatomy structures, as shown in Fig. 1. Based on it, we extend two widely used datasets, MIMIC-CXR [10] and CT-RATE [6], to create two large-scale image retrieval datasets, MIMIC-IR and CTRATE-IR, with detailed annotations of image-image similarity ordering mined from dense report annotation, serving for both training and evaluation.
@@
In summary, our contributions are threefold: (i) We propose a novel, fully automated pipeline to structure radiology reports and bridge multi-grained image-image relevance in a scalable manner. (ii) We develop MIMIC-IR and CTRATE-IR, two large-scale and comprehensive datasets accompanied by evaluation benchmarks for Chest X-ray and Chest CT image retrieval, with detailed annotations capturing image-image similarity ordering based on regional findings. (iii) We present two state-of-the-art image retrieval systems, RadIR-CXR and RadIR-ChestCT, which demonstrate superior performance in global image retrieval and achieve substantial advancements in image retrieval conditioned on anatomies.
@@ 3 Dataset Construction
In this section, we quantify image-to-image similarity ordering by mining their paired radiology reports, as shown in Fig. 1. We first introduce the data sources in Sec. 3.1. Then, we detail the two main procedures in the proposed pipeline: report decomposition in Sec. 3.2 and relevance quantification in Sec. 3.3.
3.1 Date Source
We utilize two widely used datasets: MIMIC-CXR[10] is the largest chest X-ray dataset, containing 377,110 image-report pairs; while CT-RATE[6] is a large-scale chest CT dataset with 25,692 non-contrast CT volume-report pairs. These reports include detailed descriptions of radiological findings and impressions, which are essential for defining clinically meaningful similarities between images.
@@ 4 RadIR
In this section, we present the details to build RadIR based on the datasets we construct above. The training procedure includes two stages. In Sec. 4.1, we pre-train the CLIP-style model for unconditional image retrieval; In Sec. 4.2, we extend the pre-trained model for the retrieval task conditioned on a text query.
@@ 4.1 Unconditional Image Retrieval
As shown in Fig. 2(a), in this setting, we directly encode the raw images and reports without considering extra text queries. We adopt a typical CLIP-style [16] model with a Vision Transformer based image encoder $\Phi_{\text{visual}}(\cdot)$ and a BERT-based text encoder $\Phi_{\text{text}}(\cdot)$:
@@
where $\mathbf{T}\in\mathbb{R}^{N\times N}$ is a text-text similarity matrix calculated via RaTEScore, as illustrated in Sec. 3. $\lambda_{1},\lambda_{2},\lambda_{3}$ are hyper-parameters. $\mathcal{L}_{\text{MIL}}$ is a variant of the classic infoNCE loss:
@@ 4.2 Text-Conditioned Image Retrieval
As shown in Fig. 2(b), given a conditional query $Q$, we employ a fusion module $\Phi_{\text{fusion}}$ to extend the model for conditional image retrieval:
@@
Training Objectives. In contrast to the global similarity matrix $\mathbf{T}$ based on complete reports in Sec. 4.1, we introduce $\mathbf{T}_{Q}$ as an anatomy-conditioned similarity matrix constructed from regional findings, based on equation 3. Meanwhile, the predicted conditional image-image similarity result $\mathbf{S}_{f2f}$ is derived from the dot product of fused features. We then apply triplet loss on them:
@@ 5 Experiment Settings and Results
We validate RadIR on MIMIC-IR and CTRATE-IR, with both unconditional image retrieval, image to report retrieval, and image retrieval conditioned on anatomy name. In all experiments, we follow the official train-test split of MIMIC-CXR and CT-RATE. In this section, we first introduce our baselines in Sec. 5.1 and benchmark metrics in Sec. 5.2; Then, we analyze the experiment results in Sec. 5.3 and Sec. 5.4.
5.1 Baseline
Table 2: Conditional image retrieval results on MIMIC-IR. The results were averaged and aggregated by anatomical region for presentation. Anatomies are sorted in descending order of their frequency in train set, with the ‘head’ regions at the top and the ‘tail’ regions at the bottom. The best results for each anatomy are bolded. Greener suggests higher improvement over baselines.
@@ 5.1 Baseline
5.2 Metric
[…]
5.3 Results on Unconditional Retrieval
As demonstrated in Table 1, after fine-tuning, RadIR consistently exceeds the state-of-the-art CLIP models on image-image retrieval task, and on both CXR and Chest CT datasets. Notably, RadIR can also be applied for image-report retrieval and achieves notable improvement over baselines. These results highlight that RadIR can performs effectively in these traditional retrieval tasks.
Table 3: Conditional image retrieval results on CTRATE-IR. Anatomies are in descending order of their frequency in train set, with the ‘head’ anatomies at the top and the ‘tail’ anatomies at the bottom. The best results for each anatomy are bolded. Greener suggests higher improvement over baseline.
@@ 5.3 Results on Unconditional Retrieval
5.4 Quantitative Results on Conditional Image Retrieval
Table. 2 shows that RadIR outperforms baselines in 9 out of 10 anatomical regions on CXR images, and achieves the best performance on average; While in Table. 3, RadIR consistently outperforms CT-CLIP on metrics. In addition, we observe that RadIR performs better on tail anatomies less frequently mentioned in the report. We hypothesize that this is because baselines trained on image-text pairs exhibit a bias towards more frequent anatomies. While RadIR, supporting conditional retrieval, can effectively adapt its focus to the queried anatomy, demonstrating superior robustness and versatility.
Figure 3: Qualitative comparison of conditional image retrieval results on MIMIC-IR and CTRATE-IR. Heart, lung, pleural, liver, aorta and breast are used as query conditions respectively. The query image and the top 1 retrieved images from RadIR and baselines are presented, each with regional findings attached for reference.
5.5 Qualitative Results on Conditional Image Retrieval
We further conduct a qualitative comparison in Fig. 3. It shows that RadIR can accurately retrieve images with similar findings to the query images in specific anatomical structures. For instance, for a male case presenting with bilateral gynecomastia in CTRATE-IR (the last row), RadIR can successfully retrieve a reference case exhibiting the same condition, where as CT-CLIP incorrectly matched a female case.

### v2

@@ Abstract
Keywords:
Image Retrieval Medical Imaging Vision-Language Pre-training.
@@ 1 Introduction
The objective of this paper is to develop an image retrieval system for medical applications that ranks instances in a retrieval set based on their relevance to a query, which includes a radiology image and an optional text condition indicating the region to focus on, i.e., the name of anatomy. Such a system has broad implications in enhancing clinicians’ ability to identify similar cases, supporting diagnosis and treatment planning, and facilitating medical education and research [2, 17, 15, 5]. Furthermore, in building generalist models [7, 20, 22], retrieval-augmented generation (RAG) plays a crucial role in reducing hallucinations and supporting case-based reasoning by grounding outputs in retrieved evidence.
@@
To address the challenges we propose a novel medical image ranking pipeline by mining the multi-grained annotations from corresponding radiology reports. Specifically, given a certain anatomy structure, we first standardize the paired reports and extract the relative findings. Then, we adopt the text-level similarity ranking of the findings based on well-designed language-wise metrics [23], to, in turn, represent the image similarity ranks regarding this anatomy structure. This pipeline enables the construction of multi-granular similarity ranking training data in a scalable and automated manner, for both global image matching and fine-grained retrieval conditioned on anatomy structures, as shown in Figure 1. Based on it, we extend two widely used datasets, MIMIC-CXR [11] and CT-RATE [6], to create two large-scale image retrieval datasets, MIMIC-IR and CTRATE-IR, with detailed annotations of image-image similarity ordering mined from dense report annotation, serving for both training and evaluation.
@@
In summary, our contributions are threefold: (i) We propose a novel, fully automated pipeline to structure radiology reports and bridge multi-grained image-image relevance in a scalable manner. (ii) We develop MIMIC-IR and CTRATE-IR, two large-scale and comprehensive datasets accompanied by evaluation benchmarks for Chest X-ray and Chest CT image retrieval, with detailed annotations capturing image-image similarity ordering based on regional findings. (iii) We present two state-of-the-art image retrieval systems, RadIR-CXR and RadIR-ChestCT, which demonstrate superior performance in global image retrieval and substantial advancements in image retrieval conditioned on anatomies.
@@ 3 Dataset Construction
In this section, we propose an automatic pipeline to quantify image-to-image similarity ordering, via mining their paired radiology reports, as shown in Figure 1. We first introduce the data sources in Section 3.1. Then, we detail the two main procedures in the pipeline: report decomposition in Section 3.2 and relevance quantification in Section 3.3.
3.1 Data Sources
We utilize two widely used datasets: MIMIC-CXR [11] is the largest chest X-ray dataset, containing 377,110 image-report pairs; while CT-RATE [6] is a large-scale chest CT dataset with 25,692 non-contrast CT volume-report pairs. These reports include detailed radiological findings and impressions, which are essential for defining clinically meaningful similarities between images.
@@ 4 RadIR
In this section, we present the details to build RadIR based on the datasets we construct above. The training procedure includes two stages. In Section 4.1, we pre-train the CLIP-style model for unconditional image retrieval; In Section 4.2, we extend the pre-trained model for retrieval task conditioned on a text query.
@@ 4.1 Unconditional Image Retrieval
As shown in Figure 2(a), in this setting, we directly encode the raw images and reports without considering extra text queries. We adopt a typical CLIP-style [18] model with a Vision Transformer based image encoder $\Phi_{\text{visual}}(\cdot)$ and a BERT-based text encoder $\Phi_{\text{text}}(\cdot)$:
@@
where $\mathbf{T}\in\mathbb{R}^{N\times N}$ is a text-text similarity matrix calculated via RaTEScore, as illustrated in Section 3. $\lambda_{1},\lambda_{2},\lambda_{3}$ are hyper-parameters. $\mathcal{L}_{\text{MIL}}$ is a variant of the classic infoNCE loss:
@@ 4.2 Text-Conditioned Image Retrieval
As shown in Figure 2(b), given a conditional query $Q$, we employ a fusion module $\Phi_{\text{fusion}}$ to extend the model for conditional image retrieval:
@@
Training Objectives. In contrast to the global similarity matrix $\mathbf{T}$ based on complete reports in Section 4.1, we introduce $\mathbf{T}_{Q}$ as an anatomy-conditioned similarity matrix constructed from regional findings, based on equation 3. Meanwhile, the predicted conditional image-image similarity result $\mathbf{S}_{f2f}$ is derived from the dot product of fused features. We then apply triplet loss on them:
@@ 5 Experiment Settings and Results
Table 2: Conditional image retrieval results on MIMIC-IR. Recall scores are averaged and aggregated by anatomical region, and presented in percentage. Anatomies are sorted in descending order of their frequency in train set, with the ‘head’ regions at the top and the ‘tail’ regions at the bottom. The best results for each anatomy are bolded. Greener suggests higher improvement over baselines.
@@
Table 3: Conditional image retrieval results on CTRATE-IR. Recall scores are in percentage. Anatomies are in descending order of their frequency in train set, with the ‘head’ anatomies at the top and the ‘tail’ anatomies at the bottom. The best results for each anatomy are bolded. Greener suggests higher improvement over baseline.
@@
We validate RadIR on MIMIC-IR and CTRATE-IR, with both unconditional image retrieval, image to report retrieval, and image retrieval conditioned on anatomy name. In all experiments, we follow the official train-test split of MIMIC-CXR and CT-RATE. In this section, we first introduce our baselines in Section 5.1 and benchmark metrics in Section 5.2; Then, we analyze the experiment results in Section 5.3 and Section 5.4.
5.1 Baseline
5.2 Metric
[…]
5.3 Results on Unconditional Retrieval
As demonstrated in Table 1, after fine-tuning, RadIR consistently exceeds the state-of-the-art CLIP models on image-image retrieval task, and on both CXR and Chest CT datasets. Notably, RadIR can also be applied for image-report retrieval and achieves notable improvement over baselines. These results highlight that RadIR can perform effectively in these traditional retrieval tasks.
5.4 Results on Conditional Image Retrieval
Table 2 shows that RadIR outperforms baselines in 9 out of 10 anatomical regions on CXR images, and achieves the best performance on average; While in Table 3, RadIR consistently outperforms CT-CLIP on metrics. In addition, we observe that RadIR performs better on tail anatomies less frequently mentioned in the report. We hypothesize that this is because baselines trained on image-text pairs exhibit a bias towards more frequent anatomies. While RadIR, supporting conditional retrieval, can effectively adapt its focus to the queried anatomy, demonstrating superior robustness and versatility.
@@ 6 Conclusion
{credits}
6.0.1 Acknowledgements
This study was funded by National Key R&D Program of China (No. 2022ZD0161400).
[…]
The authors have no competing interests to declare that are relevant to the content of this article.
Appendix 0.A Appendix
Figure 3: Qualitative comparison of conditional image retrieval results on MIMIC-IR and CTRATE-IR. Heart, lung, pleural, liver, aorta and breast are used as query conditions respectively. The query image and the top 1 retrieved images from RadIR and the strongest baselines are presented, each with regional findings attached for reference.